Shift report 20000724 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Mon Jul 24 2000 - 08:06:59 EDT

  • Next message: Konstantin Olchanski: "Re: Shift report 20000724 00:00-08:00"

    ------------------------------------------------------------
    Date: Mon Jul 24 08:06:59 2000, Shift: 00:00-08:00
    Supervisor: PC, with: PS
    Log book#3 pages 61 - 65
    
    Shift Summary:
    ------------------------------------------------------------
    BEAM :
    There has been beam at 70 GeV almost all night with a high
    number of ions in both beam.
    1 -  5 am, 915 / 353, run 1908, ~1835 good / 40 k
    5:30 - ?, 1051 /  595!, run 1909, ~1300 trigger 6 / 28000
    Looking at the H1 and H2 time distributions there is a band 
    across the slats, and for TOFW there are 2 bands seperated 
    by 1000 channels =  25 ns ? (plot on page 65)
    
    PROBLEMS :
    The terminal that show the scaler plots / beam status does
    not respond to the mouse, keyboard or turning the monitor
    on/off, but stays in screensave mode. We called and left a 
    message to ITD. ( This happened at 2:30 )
    SuperMon is running very slowly ( probably because of all 
    the new TPC plots ). It died a few times for us. Errors are
    noted in log book at page 63. Better histograms are needed
    for all detectors.
    We tried to look for tracks in T2 but found none. Improved
    monitor with cuts on triggers is really needed! Tracks
    in all the other TPCs.
    
    QUESTIONS :
    Should fb1, fb3 be included in the fb & camac pedestal
    runs ?
    Should the anode voltage be on the TPCs when we are doing
    pedstal runs for the TPCs ?                                                                                                                                                                    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Mon Jul 24 2000 - 08:10:40 EDT